OpenStudy (anonymous):

how do you graph –4x – 2y = 8

OpenStudy (anonymous):

–4x – 2y = 8 Solve for y :)

OpenStudy (anonymous):

-2y=8+4x add 4x to both sides y= -4-2x divide by -2 both sides